OpenStudy (anonymous):

Abel was wrapping gifts for an upcoming party. He began with 20 1/2 feet of paper. He used 5 5/12 feet on the first gift and 3 1/3 feet on the second gift. If Abel bought two additional rolls of paper that each contained 10 1/2 feet, how many feet of paper does he now have left for wrapping?

OpenStudy (anonymous):

@openstudier101x

OpenStudy (anonymous):

@Panda.exe

OpenStudy (panda.exe):

i will help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well first you subtract what he used from the initial amount he had, and then add how much additional paper he bought. To do so, you need all fractions to have the same denominator. You already have one fraction with a denominator of 12, and the other fractions all have denominators that are factors of 12, so “enlarge” each fraction to have a denominator of 12: 1/2 = 6/12 1/3= 4/12

OpenStudy (anonymous):

Now you can add and subtract by using these fractions instead of the given ones. Subtract first: 20 6/12 - 5 5/12 = 15 1/12, so after wrapping the first gift he has 15 1/12 paper left. Now for the second gift: 15 1/12- 3 4/12 = 12 9/12, so he now has 12 and 9/12 paper left after wrapping the first and second gifts. Here comes the easy part. Add this remaining amount of paper to the amount of paper he bought. He had 2 additional roles rolls of paper that each contained 10 1/2, or 10 6/12 feet, so that gives us: 12 9/12 + 10 6/12= 23 3/12, so we have 23 3/12 paper after buying the first roll, now on the the second roll... 23 3/12 + 10 6/12 = 33 9/12...But WAIT this is not the final answer yet; we have to simplify 9/12 now: 9/12 = 3/4. So, Abel now has 33 and 3/4 paper left for wrapping.

OpenStudy (anonymous):

@AwesomeSays3!

OpenStudy (anonymous):

So, your final answer is that Abel has 33.75 feet of paper left for wrapping, or about 34 feet of paper left for wrapping.

OpenStudy (anonymous):

Happy studying! :)